OpenStudy (anonymous):

the length of a new rectangle playing field is 3 yards longer than quadruple the width. if the perimeter of the rectangle playing field is 476 yards, what are its dimensions ?

OpenStudy (anonymous):

2x+2y=476 x=3+4y and so 2(3+4y)+2y=476 6+10y=476 or y = 47 x=191

OpenStudy (anonymous):

y= width x= length